Liquidating Partnerships—Deficiency Prior to liquidating their partnership, Pepper and Morrison had capital accounts of $21,000 and $85,000, respectively. The partnership assets were sold for $40,000. The partnership had no liabilities. Pepper and Morrison share income and losses equally.
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount
Assets = Capital + Liabilities
Capital + Liabilities = 21000 + 85000 +0             106,000.00
Sale Value of assets               40,000.00
Loss on Sale of assets = 106000 - 40000               66,000.00
Loss to be shared equally = 66000/2               33,000.00
Particulars Pepper Morrison
Capital Accounts               21,000.00              85,000.00
Loss to be shared equally             (33,000.00)            (33,000.00)
Net Capital to be repaid             (12,000.00)              52,000.00